OpenStudy (anonymous):

Complete the square x^2+4x+?=(x+2)^2

OpenStudy (anonymous):

4

OpenStudy (anonymous):

you will see it if you square the right hand side. you get \[(x+2)^2=(x+2)(x+2)=x^2+4x+4\]